OpenStudy (anonymous):

what is the soluion o 5+x/8=-4

OpenStudy (anonymous):

x=-27

OpenStudy (anonymous):

i disagree. heres why. your first step is to multiply both sides by 8, this gets rid of your fraction so, \[(8) (5+x/8) = -4(8)\] now you have \[5+x=-32\] from there subtract 5 on both sides \[5+x - 5=-32-5\] so now youre left with \[x=-37\]

OpenStudy (anonymous):

correct answer is -72

OpenStudy (anonymous):

add 8 to both sides\[8(5+x/8)=-4(8)\] \[40+x=-32\]then make sure the number and variable are on different sides by subtracting 40 from both sides. And your answer will be:\[x=-72\]